Grade 12Physical Chemistry

Tthe rms speed of ch4 is 3.8x10 raised to 4 cm per second.The total kinetic energy of a mixture of 0.8 moles of ch4 and 0.2 moles of c4div0 is?

JRP
7 Years agoGrade 12

1 Answer

Samyak Jain
7 Years ago
We know that average kinetic energy of gas molecules is (3/2)nRT,
n = number of moles of gas, R = gas constant, T = temperature of gas.
So, kinetic energy of 0.8 mol CH4 gas is (3/2)(0.8)RT and that of 0.2 mol C4H10 is (3/2)(0.2)RT.
We need to calculate RT.
We know that rms speed of gas molecules is \sqrt{3RT/M}  =  \sqrt{3RT/16}  for CH4 gas. (M is molar mass of gas)
\sqrt{3RT/16}  =  3.8 x 102  in SI units. Square both sides.
3RT/16 = 14.44 x 104  \Rightarrow  RT = 14.44 x 104 x 16/3.
So, average kinetic energy of 0.8 mol CH4 = (3/2)(0.8)(14.44 x 104 x 16/3) J = 924160 J = 924.16 kJ
and that of 0.2 mol C4H10 = (3/2)(0.2)(14.44 x 104 x 16/3) J = 231040 J = 231.04 kJ.